Hello, can someone please help me solve this?

How much kinetic energy would an electron depart a magnesium plate, with work function of ø= 5.86x10^ -11J, if light with a wavelength of 100 nm was shined on it?
Answer #1

Solution:

Using the relation:

=> K.E = h * v - work funciton

=> K.E = h * c / wavelength - work function

=> K.E = 6.626 * 10^-34 * 3 * 10^8 / (100 * 10^-9) - 5.86 * 10^-11

=> K.E = 19.878 * 10^-19 - 5.86 * 10^-11  

if we calculate the above relation, we will get negative value for kinetic energy which is not possible

Hence, we can say electron will not eject from the plate and there will be no kinetic energy at all if electron is not getting depart from the plate.

Implies , K.E = 0 J

please comment for queries.

thanks.